OpenStudy (chriss):

Determine whether the given map \[\phi\] is an isomorphism of the first binary structure with the second. If it is not an isomorphism, why not? \[<\mathbb{Q},*>\ with\ <\mathbb{Q},*>\ where\ \phi(x)=x^{2}\ for\ x \in \mathbb{Q}\] I have the one-to-one requirement \[If\ \phi(x)=\phi(y) then\ x^{2}=y^{2}\] \[Then\ \sqrt{x^{2}}=\sqrt{y^{2}}\] So x=y Thus the map is one-to-one. Now I need to show that it is onto and I need to show \[\phi(x*y)=\phi(x)*\phi(y)\]

OpenStudy (anonymous):

i think this is not one to one

OpenStudy (anonymous):

for example \[\phi(2)=\phi(-2)\]

OpenStudy (anonymous):

and the mistake is in this line \[x^2=y^2\iff x=y\] should be \[x^2=y^2\iff x = \pm y\]

OpenStudy (anonymous):

also it is not onto since nothing gets mapped to negative rationals

OpenStudy (chriss):

right... I see I was totally off on this now. Thanks!

OpenStudy (anonymous):

yw

OpenStudy (chriss):

Hopefully you won't mind verifying one more for me... Same question but different binary structures... I think I have the answer but would like to run it by you if I could.. for this one the binary structure is \[<M_{2}(\mathbb{R}),*>\ with <\mathbb{R},*> where\ \phi(A)\ is\ the\ determinant\ of\ matrix\ A\] This one is also not an isomorphism because multiplication of real numbers is commutative, where Matrix multiplication is not.

OpenStudy (anonymous):

no it is not an isomorphism because the set of two by two matrices is the same as \[\mathbb{R^4}\]

OpenStudy (anonymous):

it is a homomorphism however, because the determinant of the product is the product of the determinants

OpenStudy (anonymous):

that is \[\phi(AB)=\phi(A)\phi(B)\]

OpenStudy (anonymous):

i am assuming the " * " in both cases means multiplication yes?

OpenStudy (anonymous):

in any case it is certainly not injective because an infinite number of matrices can have the same determinant

OpenStudy (chriss):

awesome, thanks again. We haven't got to homomorphisms yet, but I will note this for when we do. Yes, I didn't see a symbol for using a dot for multiplication. I know it's confusing since * can be a generic binary operation symbol too but I thought it would make sense in the context of the question. I do appreciate your help though. I'll add a medal to one of the other questions you have answered since I can't give you two here lol. Thanks a bunch :)

OpenStudy (anonymous):

no problem, i have medals to spare ( and i can't seem to redeem them at the package store)